# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ 4.3 Existence of a single-lobe kink state
$-\frac{c_{1}^{2}}{2}[\phi_{1,\beta}^{\prime}(L)]^{2}=\cos(\phi_{1,\beta}(L))-\cos(\phi_{1,\beta}(0))=-2\beta^{2}\mathrm{cn}^{2}\Big{(}\frac{L}{c_{1}}+K(\beta);\beta\Big{)}=0.$ (4.22)
Therefore, $\phi_{1,\beta}^{\prime}(L)=0$, which is clearly not possible. We note from the profile of the elliptic function $\operatorname{sn}(\cdot)$ that $-1\leqq\operatorname{sn}(\frac{x}{c_{1}}+K)\leqq 0$ for $x\in(b,L]$ and for $x=L$, $K<\frac{L}{c_{1}}<2K$ and hence $\operatorname{sn}(\frac{L}{c_{1}})>0$ and $\mathrm{cn}(\frac{L}{c_{1}})<0$.
Next, to show the continuity condition $\phi_{1,k}(L)=\phi_{2,a}(L)$ for $a=a(k)>0$ we shall follow the same strategy as in the proof of Proposition 4.7. Indeed, we get the first same expression for the shift $a$ as in (4.9) by considering again $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ but with $\beta\in(0,\frac{\pi}{4}]$. Thus, from relation $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ and $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Thus, we obtain (4.9) with $a>0$ and a similar relation as in (4.11). Next, we determine the right $k\in(0,k_{0})$ such that $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$. Initially we get
@@
Therefore, $\lim_{k\to 0}a(k)=0$ and $\lim_{k\to k_{0}}a(k)=0$ ($a(k_{0})=0$). We note that, compared to the negative case of $a$, $a=a(k)$ is no longer a strictly increasing function of $k$. In Figure 15 we show a generic profile of $a(k)$ given by formula in (4.13) and we notice that there is always a unique value $k_{\mathrm{crit}}\in(0,k_{0})$ such that $a^{\prime}(k_{\mathrm{crit}})=0$.
@@
Now, as relations for $\phi_{1,k}^{\prime}(L)$ and $\phi_{2,a}^{\prime}(L)$ in (4.14) and (4.15), respectively, holds true, we obtain $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$ if and only if $k$ satisfies the following relation
@@
Let $Z<0$, then from equation $H(k)=Z$ in (4.18) with $H$ defined in (4.17) for $k\in(0,k_{0})$, we need to have a priori that $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)$, for some $k\in(0,k_{0})$. Therefore, for $\frac{c_{1}}{2c_{2}}\geqq k_{0}$ there is not a single-lobe kink state. This finishes the proof. ∎
@@
    Let us illustrate the case $L=3\pi$, $c_{1}=1$, with $k_{0}$ satisfying $K(k_{0})=3\pi$ ($k_{0}\approx 1$). Figure 16 shows the profile of $F(k)=k\operatorname{sn}\big{(}K(k_{0});k\big{)}$ for $k\in(0,k_{0})$. Now, $F(k)=0$ if and only if $r_{1}\approx 0.84$ and $r_{2}\approx 0.9987$ (see Figures 17 and 17). By chosing $c_{2}=2$, we have that $F(k)=\frac{1}{4}$ if and only if $k_{1}\approx 0.32$, $k_{2}\approx 0.81$ and $k_{3}\approx 0.999$. But the right $k_{i}$ that produces a single-lobe kink profile ($\phi_{1,k}^{\prime}(x)<0$ for all $x\in[0,3\pi]$) will be exactly $k_{3}\approx 0.999$. In Figures 18, 18 and 18 we show the profile of $\phi_{1,k_{i}}(x)$ in (4.8) with $x\in[K(k_{i}),3\pi]$.
@@
    For the case $Z<0$ and $c_{2}$ such that $k_{0}>\frac{c_{1}}{2c_{2}}$, some technical problems arise in studying equation $H(k)=Z$ in (4.18) to obtain a general picture of existence of single-lobe kink profile as was that obtained in Proposition 4.7. The problem here is that the mapping $k\to k\operatorname{sn}(L/c_{1};k)$, $k\in(0,k_{0})$, may have a profile with several oscillations that do not allow us to clearly study the mapping $H$ in (4.17). A similar situation is obtained for $Z>0$.
@@ 5 Spectral structure associated to positive single-lobe kink states
In what follows, we establish our main strategy in order to study the eigenvalue problem in (5.2). More precisely, we will reduce the eigenvalue problem for $(\mathcal{L},D_{Z,0})$ into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on $[-L,L]$, and one for the operator $\mathcal{L}_{a}$ with $\delta$-type boundary conditions on the half-line $(0,\infty)$. Thus, we employ the so-called splitting eigenvalue method on a tadpole graph (see Angulo [6, 7]).
@@
Suppose that $(f,g)\in D_{Z,0}$ with $g(0)\neq 0$, satisfies the relation $\mathcal{L}(f,g)^{\top}=\lambda(f,g)^{\top}$ for $\lambda\leqq 0$. Then, we obtain the following two eigenvalues problems
@@
$f(-L)=f(L),\;\;f^{\prime}(L)-f^{\prime}(-L)=\Big{[}\frac{g^{\prime}(0+)}{g(0+)}+Z\Big{]}f(L)\equiv\theta f(L).$
Now, we consider the following eigenvalue problem for $\mathcal{L}_{1}$ with real coupled self-adjoint boundary conditions determined by $\alpha\in\mathbb{R}$,
$(RC_{\alpha}):\;\;\left\{\begin{array}[]{ll}\mathcal{L}_{1}y(x)=\eta y(x),\quad\quad x\in(-L,L),\\ y(L)=y(-L),\\ y^{\prime}(L)-y^{\prime}(-L)=\alpha y(L).\end{array}\right.$ (5.7)
Thus, $f$ in (5.6) satisfies the real coupled problem $(RC_{\alpha})$ in (5.7) wth $\alpha=\theta$ and $\eta=\lambda$. We shall verify that $\theta=0$, which proves the Lemma. Indeed, we consider $K_{\theta}=[k_{ij}]$ the $2\times 2$-matrix associated to (5.7) given by $k_{11}=1$, $k_{12}=0$, $k_{21}=\theta$ and $k_{22}=1$ ($\text{det}(K_{\theta})=1$), and by $\eta_{n}=\eta_{n}(K_{\theta})$, $n\in\mathbb{N}_{0}$, the eigenvalues for the $(RC_{\theta})$-problem in (5.7). We also consider $\mu_{n}=\mu_{n}(K_{\theta})$ and $\nu_{n}=\nu_{n}(K_{\theta})$, $n\in\mathbb{N}_{0}$, the eigenvalues problems in (5.7) induced by $K_{\theta}$ with the following boundary conditions
[…]
respectively. We recall that if $y_{n}$ is an eigenfunction of $\mu_{n}$, then $y_{n}$ is unique up to constant multiples and has exactly $n$ zeros in $(-L,L)$, $n\in\mathbb{N}_{0}$ (a similar result is obtained for $u_{n}$ an eigenfunction of $\nu_{n}$, $n\in\mathbb{N}_{0}$). Next, by Theorem 4.8.1 in Zettl [52] we have that $\nu_{0}$ and $\eta_{0}$ are simple and, in particular, we have the following partial distribution of eigenvalues,
$\nu_{0}\leqq\eta_{0}<\{\mu_{0},\nu_{1}\}<\eta_{1}\leqq\{\mu_{1},\nu_{2}\}\leqq\eta_{2},$ (5.9)
where the notation $\{\mu_{n},\nu_{m}\}$ is used to indicate either $\mu_{n}$ or $\nu_{m}$ (but no comparison is made between $\mu_{n}$ and $\nu_{m}$). Next, we observe that $u^{\prime}_{0}(L)=0$ and so $\theta=0$ (because by (5.9) we have $\nu_{0}<\mu_{0}$ and so $u_{0}(\pm L)\neq 0$). Indeed, since $\nu_{0}$ is simple and the profile-solution $\phi_{1}$ is even, we need to have that $u_{0}$ is even or odd, but as $u_{0}$ has not zeros in $[-L,L]$ we get that $u_{0}$ is actually even. Thus, $u^{\prime}_{0}$ is odd and $u^{\prime}_{0}(\pm L)=0$. This finishes the proof. ∎
@@ 5.1 Morse index for $\mathcal{L}_{Z}$
We note that always we have $\psi^{\prime}_{a}\neq 0$ on $[0,\infty)$. Thus for $f\in C_{0}^{\infty}(0,\infty)$ we obtain
[…]
@@
Notice that it is possible to have, in Lemma 5.2, that $n(\mathcal{L}_{a})=0$ for some $W_{Z}$-domain. Indeed, by considering $Z_{a}=-\psi^{\prime\prime}_{a}(0)/\psi^{\prime}_{a}(0)$ we have $\psi^{\prime}_{a}\in W_{Z_{a}}$ and $\mathcal{L}_{a}\psi^{\prime}_{a}=0$. Thus, as $\psi^{\prime}_{a}<0$ on $[0,\infty)$ follows from Sturm-Liouville oscillation theory on half-lines that zero is the smallest eigenvalue for $(\mathcal{L}_{a},W_{Z_{a}})$ (note that the essential spectrum for $\mathcal{L}_{a}$ is $\sigma_{\mathrm{ess}}(\mathcal{L}_{a})=[1,\infty)$).
By Theorem A.7 in Appendix we obtain that the Morse index of $(\mathcal{L},D_{Z,0})$ satisfies $n(\mathcal{L})\leqq 2$. Upon application of Lemmata 5.1 and 5.2 we obtain the following result.
@@
Consider the pair $(\mathcal{L},D_{Z,0})$. Suppose that there are $(f_{i},g_{i})\in D_{Z,0}$ with $g_{i}(0)\neq 0$, $i=1,2$, such that $\mathcal{L}(f_{i},g_{i})^{\top}=\lambda_{i}(f_{i},g_{i})^{\top}$ with $\lambda_{i}<0$. Then, $\lambda_{1}=\lambda_{2}$ and $g_{1}=\alpha g_{2}$ on $[0,\infty)$.
@@
By Theorem A.7 in Appendix, we known that $n(\mathcal{L})\leqq 2$. Suppose that $n(\mathcal{L})=2$ and consider $\lambda_{0},\lambda_{1}$ being the negative eigenvalues (counting multiplicity).
@@
        The profile $\zeta_{\lambda_{0}}$ is not identically zero. Indeed, suppose that $\zeta_{\lambda_{0}}\equiv 0$; then $\xi_{\lambda_{0}}$ satisfies
        $\left\{\begin{array}[]{ll}\mathcal{L}_{1}\xi_{\lambda_{0}}(x)=\lambda_{0}\xi_{\lambda_{0}}(x),\quad\quad x\in(-L,L),\\ \xi_{\lambda_{0}}(L)=\xi_{\lambda_{0}}(-L)=0,\\ \xi^{\prime}_{\lambda_{0}}(L)=\xi^{\prime}_{\lambda_{0}}(-L).\end{array}\right.$ (5.15)
        Next, from the Dirichlet condition (which implies that the eigenvalue $\lambda_{0}$ is simple and so $\xi_{\lambda_{0}}$ is odd or even) and from oscillations theorems of the Floquet theory (which imply that the number of zeros of $\xi_{\lambda_{0}}$ is even on $[-L,L)$), we need to have that $\xi_{\lambda_{0}}$ is odd. Then, by Sturm-Liouville theory there is an eigenvalue $\theta$ for $\mathcal{L}_{1}$, such that $\theta<\lambda_{0}$, with associated eigenfunction $\chi>0$ on $(-L,L)$, and $\chi(-L)=\chi(L)=0$. Now, let $\mathcal{Q}_{\mathrm{Dir}}$ be the quadratic form associated to $\mathcal{L}_{1}$ with Dirichlet domain, namely, $\mathcal{Q}_{\mathrm{Dir}}:H^{1}_{0}(-L,L)\to\mathbb{R}$ and
@@
        ~$\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (5.17)
@@
        for any $\gamma\in\mathbb{R}$. Then, from the extension theory for symmetric operators we have that the family (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ represents all the self-adjoint extensions of the symmetric operator $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ defined by
        $\mathcal{N}_{0}=\widetilde{\mathcal{L}},\quad D(\mathcal{N}_{0})=\{f\in H^{2}(\mathbb{R}):f(0)=0\}.$
        Moreover, the deficiency indices of $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ are given by $n_{\pm}(\mathcal{N}_{0})=1$ (see Albeverio et al. [4] and Proposition A.3 in Appendix below). Now, the even tail-profile $\psi_{\mathrm{even}}$ satisfies $\psi_{\mathrm{even}}^{\prime}(x)\neq 0$ for all $x\neq 0$, and so from the relation
@@
        we can see easily that $\langle\mathcal{N}_{0}f,f\rangle\geqq 0$ for all $f\in D(\mathcal{N}_{0})$. Then, from the extension theory (see Proposition A.5 in Appendix) we obtain that the Morse index for the family (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})$ satisfies $n(\widetilde{\mathcal{L}})\leqq 1$, for all $\gamma\in\mathbb{R}$. Therefore, since $\zeta_{\mathrm{odd}}\in D_{\delta,\gamma}$, for any $\gamma$, and ~$\widetilde{\mathcal{L}}\,\zeta_{\mathrm{odd}}=\lambda_{0}\zeta_{\mathrm{odd}}$ on $\mathbb{R}$, we have $n(\widetilde{\mathcal{L}})=1$ and $\lambda_{0}$ will be the smallest negative eigenvalue for $\widetilde{\mathcal{L}}$ on $\delta$-interactions domains. Then, by Theorem A.8 in Appendix (Perron Frobenius property for $\widetilde{\mathcal{L}}$), $\zeta_{\mathrm{odd}}$ needs to be positive which is a contradiction. Therefore, $\zeta_{\lambda_{0}}(0)\neq 0$.
@@
        $\zeta_{\lambda_{0}}$ and $\xi_{\lambda_{0}}$ are positive profiles: from $\zeta_{\lambda_{0}}(0)\neq 0$ and Lemma 5.1, the pair $(\zeta_{\lambda_{0}},\lambda_{0})$ satisfies the $(\delta BP)$-boundary problem in (5.6) and so by Lemma 5.2 we obtain $n(\mathcal{L}_{a})=1$. Therefore, by oscillation theory we can choose $\zeta_{\lambda_{0}}>0$ on $(0,\infty)$ (indeed, we can use the Perron-Frobenius property in Theorem A.8 for the unfold operator (~$(\widetilde{\mathcal{L}},D_{\delta,-Z})$ in (A.17)-(A.18) and also deduce the positive property of $\zeta_{\lambda_{0}}$). Next, we study $\xi_{\lambda_{0}}$. From Lemma 5.1, the pair $(\xi_{\lambda_{0}},\lambda_{0})$ satisfies the $(PBP)$-boundary problem in (5.6). We denote by $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), then $\lambda_{0}\geqq\eta_{0}$. We are going to show that $\lambda_{0}=\eta_{0}$. Indeed, we consider the quadratic form, $\mathcal{Q}_{\mathrm{per}}$, associated to the (PBP)-problem in (5.6), namely, for $h\in H_{\mathrm{per}}^{1}(-L,L)$ (with $h(L)=h(-L),h^{\prime}(L)=h^{\prime}(-L)$) we have,
[…]
        Next, for $h$ fixed as above, define $\chi=\nu\zeta_{\lambda_{0}}$ with $\nu\in\mathbb{R}$ being chosen such that $\chi(0)=\nu\zeta_{\lambda_{0}}(0)=h(L)$. Then, $(h,\chi)\in D(\mathcal{Q}_{Z})$ in (5.14). Now, by using that $\mathcal{L}_{a}\zeta_{\lambda_{0}}=\lambda_{0}\zeta_{\lambda_{0}}$ on $(0,\infty)$, we obtain
        $\mathcal{Q}_{\mathrm{per}}(h)$ $=\mathcal{Q}_{Z}(h,\chi)-\int_{0}^{\infty}(\chi^{\prime})^{2}+W_{\psi_{a}}\chi^{2}dx+Z|\chi(0)|^{2}$ (5.21)
        $=\mathcal{Q}_{Z}(h,\chi)-\langle\mathcal{L}_{a}\chi,\chi\rangle=\mathcal{Q}_{Z}(h,\chi)-\lambda_{0}\|\chi\|^{2}$
@@
        Then, $\eta_{0}\geqq\lambda_{0}$ and so $\eta_{0}=\lambda_{0}$. Therefore, from Floquet theory follows that we can choose $\xi_{\lambda_{0}}$ positive on $[-L,L]$. Even more, since $V_{\phi_{1}}$ is even, we obtain that $\xi_{\lambda_{0}}$ is even.

### v2

@@ 4.3 Existence of a single-lobe kink state
$\frac{c_{1}^{2}}{2}[\phi_{1,\beta}^{\prime}(L)]^{2}=\cos(\phi_{1,\beta}(L))-\cos(\phi_{1,\beta}(0))=-2\beta^{2}\mathrm{cn}^{2}\Big(\frac{L}{c_{1}}+K(\beta);\beta\Big)=0.$ (4.14)
Therefore, $\phi_{1,\beta}^{\prime}(L)=0$, which is clearly not possible. We note from the profile of the elliptic function $\operatorname{sn}(\cdot)$ that $1\leqq\operatorname{sn}(\frac{x}{c_{1}}+K)\leqq 0$ for $x\in(b,L]$ and for $x=L$, $K<\frac{L}{c_{1}}<2K$ and hence $\operatorname{sn}(\frac{L}{c_{1}})>0$ and $\mathrm{cn}(\frac{L}{c_{1}})<0$.
Next, to show the continuity condition $\phi_{1,k}(L)=\phi_{2,a}(L)$ for $a=a(k)>0$ we shall follow the same strategy as in the proof of Lemma 4.9. Indeed, we get the first same expression for the shift $a$ as in (4.11) by considering again $\beta\equiv\arctan(e^{-\frac{a}{c_{2}}})$ but with $\beta\in(0,\frac{\pi}{4}]$. Thus, from relation $k^{2}\operatorname{sn}^{2}(\frac{L}{c_{1}}+K;k)=\cos^{2}(2\beta)$ we get $\cos(2\beta)=-k\operatorname{sn}(\frac{L}{c_{1}}+K;k)$ and $\sin(2\beta)=\mathrm{dn}(\frac{L}{c_{1}}+K;k)$. Thus, we obtain (4.11) with $a>0$ and a similar relation as in (4.13). Next, we determine the right $k\in(0,k_{0})$ such that $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$. Initially we get
@@
Therefore, $\lim_{k\to 0}a(k)=0$ and $\lim_{k\to k_{0}}a(k)=0$ ($a(k_{0})=0$). We note that, compared to the negative case of $a$, $a=a(k)$ is no longer a strictly increasing function of $k$. In Figure 11 we show a generic profile of $a(k)$ given by formula in (A.6) and we notice that there is always a unique value $k_{\mathrm{crit}}\in(0,k_{0})$ such that $a^{\prime}(k_{\mathrm{crit}})=0$.
@@
Now, as relations for $\phi_{1,k}^{\prime}(L)$ and $\phi_{2,a}^{\prime}(L)$ in (A.7) and (A.8), respectively, holds true, we obtain $2\phi_{1,k}^{\prime}(L)=\phi^{\prime}_{2,a(k)}(L)$ if and only if $k$ satisfies the following relation
@@
Let $Z<0$, then from equation $H(k)=Z$ in (A.11) with $H$ defined in (A.10) for $k\in(0,k_{0})$, we need to have a priori that $\frac{c_{1}}{2c_{2}}<k\operatorname{sn}(L/c_{1};k)$, for some $k\in(0,k_{0})$. Therefore, for $\frac{c_{1}}{2c_{2}}\geqq k_{0}$ there is not a single-lobe kink state. This finishes the proof. ∎
@@
    Let us illustrate the case $L=3\pi$, $c_{1}=1$, with $k_{0}$ satisfying $K(k_{0})=3\pi$ ($k_{0}\approx 1$). Figure 12 shows the profile of $F(k)=k\operatorname{sn}\big(K(k_{0});k\big)$ for $k\in(0,k_{0})$.
@@
    Now, $F(k)=0$ if and only if $r_{1}\approx 0.84$ and $r_{2}\approx 0.9987$ (see Figures 13 and 13). By chosing $c_{2}=2$, we have that $F(k)=\frac{1}{4}$ if and only if $k_{1}\approx 0.32$, $k_{2}\approx 0.81$ and $k_{3}\approx 0.999$. But the right $k_{i}$ that produces a single-lobe kink profile ($\phi_{1,k}^{\prime}(x)<0$ for all $x\in[0,3\pi]$) will be exactly $k_{3}\approx 0.999$.
@@
    For the case $Z<0$ and $c_{2}$ such that $k_{0}>\frac{c_{1}}{2c_{2}}$, some technical problems arise in studying equation $H(k)=Z$ in (A.11) to obtain a general picture of existence of single-lobe kink profile as was that obtained in Proposition 4.7. The problem here is that the mapping $k\to k\operatorname{sn}(L/c_{1};k)$, $k\in(0,k_{0})$, may have a profile with several oscillations that do not allow us to clearly study the mapping $H$ in (A.10). A similar situation is obtained for $Z>0$.
@@ 5 Spectral structure associated to positive single-lobe kink states
In what follows, we establish our main strategy in order to study the eigenvalue problem in (5.2). More precisely, we will reduce the eigenvalue problem for $(\mathcal{L},D_{Z,0})$ into two classes of eigenvalue problems, one for $\mathcal{L}_{1}$ with periodic boundary conditions on [L,L][-L,L], and one for the operator $\mathcal{L}_{a}$ with $\delta$-type boundary conditions on the half-line $(0,\infty)$. Thus, we employ the so-called splitting eigenvalue method on a tadpole graph (see Angulo [5, 6]).
@@
Suppose that $(f,g)\in D_{Z,0}$ with $g(0)\neq 0$, satisfies the relation $\mathcal{L}(f,g)^{\top}=\lambda(f,g)^{\top}$ for $\lambda<0$. Then, we obtain the following two eigenvalues problems
@@
$f(-L)=f(L),\;\;f^{\prime}(L)-f^{\prime}(-L)=\Big[\frac{g^{\prime}(0+)}{g(0+)}+Z\Big]f(L)\equiv\theta f(L),$
and so $f$ satisfies the following real coupled problem
$\left\{\begin{array}[]{ll}\mathcal{L}_{1}f(x)=\lambda f(x),\quad\quad x\in(-L,L),\\ f(L)=f(-L),\\ f^{\prime}(L)-f^{\prime}(-L)=\theta f(L).\end{array}\right$. (5.7)
Then, by using oscillation theory for real coupled problems (see Theorem 4.8.1 in Zettl [41]), $\phi_{1}$ has an even single-lobe profile, Sturm comparison theorem, and $\lambda<0$, we get $\theta=0$. This finishes the proof.
@@ 5.1 Morse index for $\mathcal{L}_{Z}$
We note that always we have $\psi^{\prime}_{a}\neq 0$ on $0,\infty)$. Thus for $f\in C_{0}^{\infty}(0,\infty)$ we obtain
$\begin{split}\langle\mathcal{M}f,f\rangle&=\int_{0}^{\infty}c_{2}^{2}(\psi^{\prime}_{a})^{2}\Big|\frac{d}{dx}\Big(\frac{f}{\psi^{\prime}_{a}}\Big)\Big|^{2}dx+c_{2}^{2}f(0)\Big[\frac{f^{\prime}(0)\psi^{\prime}_{a}(0)-f(0)\psi^{\prime}_{a}(0)}{\psi^{\prime}_{a}(0)}\Big].\end{split}$ (5.10)
@@
Notice that it is possible to have, in Lemma 5.2, that $n(\mathcal{L}_{a})=0$ for some $W_{Z}$-domain. Indeed, by considering $Z_{a}=-\psi^{\prime\prime}_{a}(0)/\psi^{\prime}_{a}(0)$ we have $\psi^{\prime}_{a}\in W_{Z_{a}}$ and $\mathcal{L}_{a}\psi^{\prime}_{a}=0$. Thus, as $\psi^{\prime}_{a}<0$ on $0,\infty)$ follows from Sturm-Liouville oscillation theory on half-lines that zero is the smallest eigenvalue for $(\mathcal{L}_{a},W_{Z_{a}})$ (note that the essential spectrum for $\mathcal{L}_{a}$ is $\sigma_{\mathrm{ess}}(\mathcal{L}_{a})=[1/c_{2}^{2},\infty)$).
Upon application of Lemmata 5.1 and 5.2 we obtain the following result.
@@
Consider the pair $(\mathcal{L},D_{Z,0})$. Suppose that there are $(f_{i},g_{i})\in D_{Z,0}$ with $g_{i}(0)\neq 0$, $i=1,2$, such that $\mathcal{L}(f_{i},g_{i})^{\top}=\lambda_{i}(f_{i},g_{i})^{\top}$ with $\lambda_{i}<0$. Then, $\lambda_{1}=\lambda_{2}$ and $g_{1}=\alpha g_{2}$ on $0,\infty)$.
@@
The proof follows the strategy as in the proof of Theorems 3.1 and 3.2 in [5]. So, by the sake of completeness, we highlight the main points of the analysis. Indeed, without loss of generality, suppose $n(\mathcal{L})=2$ and consider $\lambda_{0},\lambda_{1}$ being the negative eigenvalues (counting multiplicities).
@@
        The profile $\zeta_{\lambda_{0}}$ is not identically zero: indeed, suppose that $\zeta_{\lambda_{0}}\equiv 0$; then $\xi_{\lambda_{0}}$ satisfies $\mathcal{L}_{1}\xi_{\lambda_{0}}(x)=\lambda_{0}\xi_{\lambda_{0}}(x)$, $x\in(-L,L)$, $\xi_{\lambda_{0}}(L)=\xi_{\lambda_{0}}(-L)=0$ and $\xi^{\prime}_{\lambda_{0}}(L)=\xi^{\prime}_{\lambda_{0}}(-L)$, and so from the Dirichlet condition and Oscillation Theorem of the Floquet theory, we need to have that $\xi_{\lambda_{0}}$ is odd. Then, by Sturm-Liouville theory there is an eigenvalue $\theta$ for $\mathcal{L}_{1}$, such that $\theta<\lambda_{0}$, with associated eigenfunction $\chi>0$ on (L,L)(-L,L), and $\chi(-L)=\chi(L)=0$. Now, let $\mathcal{Q}_{\mathrm{Dir}}$ be the quadratic form associated to $\mathcal{L}_{1}$ with Dirichlet domain, namely, $\mathcal{Q}_{\mathrm{Dir}}:H^{1}_{0}(-L,L)\to\mathbb{R}$ and
@@
        $\widetilde{\mathcal{L}}=-c_{2}^{2}\partial_{x}^{2}+\cos(\psi_{\mathrm{even}}),$ (5.14)
@@
        for any $\gamma\in\mathbb{R}$. Then, from the extension theory for symmetric operators we have that the family $(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ represents all the self-adjoint extensions of the symmetric operator $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ defined by $\mathcal{N}_{0}=\widetilde{\mathcal{L}}$, $D(\mathcal{N}_{0})=\{f\in H^{2}(\mathbb{R}):f(0)=0\}$, because the deficiency indices of $(\mathcal{N}_{0},D(\mathcal{N}_{0}))$ are given by $n_{\pm}(\mathcal{N}_{0})=1$ (see Albeverio et al. [4] and Proposition A.3 in Appendix below). Now, the even tail-profile $\psi_{\mathrm{even}}$ satisfies $\psi_{\mathrm{even}}^{\prime}(x)\neq 0$ for all $x\neq 0$, and so from the relation
@@
        we can see easily that $\langle\mathcal{N}_{0}f,f\rangle\geqq 0$ for all $f\in D(\mathcal{N}_{0})$. Then, from Proposition A.4 in Appendix, we obtain that the Morse index for the family $(\widetilde{\mathcal{L}},D_{\delta,\gamma})$ satisfies $n(\widetilde{\mathcal{L}})\leqq 1$, for all $\gamma\in\mathbb{R}$. Next, since $\zeta_{\mathrm{odd}}\in D_{\delta,\gamma}$, for any $\gamma$, and $\widetilde{\mathcal{L}}\,\zeta_{\mathrm{odd}}=\lambda_{0}\zeta_{\mathrm{odd}}$ on $\mathbb{R}$, we have $n(\widetilde{\mathcal{L}})=1$ and $\lambda_{0}$ will be the smallest negative eigenvalue for $\widetilde{\mathcal{L}}$ on $\delta$-interactions domains. Then, by Theorem A.5 in Appendix (Perron Frobenius property for $\widetilde{\mathcal{L}}$), $\zeta_{\mathrm{odd}}$ needs to be positive which is a contradiction. Therefore, $\zeta_{\lambda_{0}}(0)\neq 0$.
@@
        $\zeta_{\lambda_{0}}$ and $\xi_{\lambda_{0}}$ are positive profiles: from $\zeta_{\lambda_{0}}(0)\neq 0$ and Lemma 5.1, the pair $(\zeta_{\lambda_{0}},\lambda_{0})$ satisfies the $(\delta BP)$-boundary problem in (5.6) and so by Lemma 5.2 we obtain $n(\mathcal{L}_{a})=1$. Therefore, by oscillation theory we can choose $\zeta_{\lambda_{0}}>0$ on $(0,\infty)$ (indeed, we can use the Perron-Frobenius property in Theorem A.5 for the unfold operator $(\widetilde{\mathcal{L}},D_{\delta,-Z})$ in (A.3)-(A.4) and also deduce the positive property of $\zeta_{\lambda_{0}}$). Next, we study $\xi_{\lambda_{0}}$. From Lemma 5.1, the pair $(\xi_{\lambda_{0}},\lambda_{0})$ satisfies the $(PBP)$-boundary problem in (5.6). We denote by $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), then $\lambda_{0}\geqq\eta_{0}$. We are going to show that $\lambda_{0}=\eta_{0}$. Indeed, we consider the quadratic form, $\mathcal{Q}_{\mathrm{per}}:H_{\mathrm{per}}^{1}(-L,L)\to\mathbb{R}$,
        $\mathcal{Q}_{\mathrm{per}}(h)=\int_{-L}^{L}\big((h^{\prime})^{2}+V_{\phi_{1}}h^{2}\big)\,dx.$ (5.17)
        Next, for $h\in H_{\mathrm{per}}^{1}(-L,L)$ fixed, define $\chi=\nu\zeta_{\lambda_{0}}$ with $\nu\in\mathbb{R}$ being chosen such that $\chi(0)=\nu\zeta_{\lambda_{0}}(0)=h(L)$. Then, $(h,\chi)\in D(\mathcal{Q}_{Z})$ in (5.12). Now, by using that $\mathcal{L}_{a}\zeta_{\lambda_{0}}=\lambda_{0}\zeta_{\lambda_{0}}$ on $(0,\infty)$, we get immediately that
        $\mathcal{Q}_{\mathrm{per}}(h)$ $=\mathcal{Q}_{Z}(h,\chi)-\langle\mathcal{L}_{a}\chi,\chi\rangle=\mathcal{Q}_{Z}(h,\chi)-\lambda_{0}\|\chi\|^{2}$ (5.18)
@@
        Then, $\eta_{0}\geqq\lambda_{0}$ and so $\eta_{0}=\lambda_{0}$. Therefore, from Floquet theory follows that we can choose $\xi_{\lambda_{0}}$ positive on [L,L][-L,L]. Even more, since $V_{\phi_{1}}$ is even, we obtain that $\xi_{\lambda_{0}}$ is even.